OpenStudy (anonymous):

3^(2x) + 3^(x) - 2 = 0 I know that the answer is 0, but I'm having trouble solving for x.

jimthompson5910 (jim_thompson5910):

3^(2x) + 3^(x) - 2 = 0 3^(x*2) + 3^(x) - 2 = 0 (3^x)^2 + 3^(x) - 2 = 0 (3^x)^2 + (3^x) - 2 = 0 (z)^2 + (z) - 2 = 0 z^2 + z - 2 = 0 solve that for z then use the solutions of z to get the solutions of x